# Test Time Learning for Time Series Forecasting

### v1

@@ 1 Introduction
  • Hidden Layer Architectures: The ablation studies revealed that while convolutional architectures added to the TTT modules provided some improvements, Conv Stack 5 consistently delivered the best results among the convolutional variants. However, simpler architectures like Conv 3 often performed comparably, showing that increased architectural complexity did not always lead to significantly better performance.
  • Adaptability to Long-Term Predictions: The TTT-based models excelled in long-term prediction tasks, especially for really high prediction lengths like 2880. TTT-based models also excelled on increased sequence lengths as high as 5760 which is the maximum sequence length allowed by the benchmark datasets. This verified the theoretically expected supremacy of TTT based models relative to the mamba/transformer based SOTA models.
@@
In this work, we explore the potential of TTT modules in Long-Term Series Forecasting (LTSF) by integrating them into novel model configurations to surpass the current state-of-the-art (SOTA) models.
Our key contributions are as follows:
@@
  • We conduct an ablation study to assess the performance of various hidden layer architectures within our model. By testing five different convolutional configurations, we quantify their impact on model performance and provide insights into how they compare with the SOTA model.
@@ 2 Related Work
Several Transformer-based models have significantly advanced long-term time series forecasting (LTSF), with iTransformer [18] and PatchTST [21] being notable examples. iTransformer uses multimodal interactive attention to model both temporal and inter-modal dependencies, making it ideal for multivariate time series. However, its reliance on multimodal data can raise computational costs when such interactions are absent. PatchTST, inspired by Vision Transformers (ViTs) [6], divides input sequences into patches to capture long- and short-term dependencies, but its performance depends on choosing the correct patch size, and it may reduce interpretability.
Other influential models include Informer [31], which reduces complexity with sparse self-attention but can miss finer details in multivariate data; Autoformer [28], which captures periodicity and trends but struggles with non-periodic data; Pyraformer [16], which handles multi-scale patterns via a hierarchical structure but increases computational cost; and Fedformer [32], which combines time and frequency representations to reduce overhead but may underperform on noisy time series. Each model improves LTSF in specific ways but also introduces its own limitations.
@@
[8, 9, 10] are efficient sequence models designed for long-term time series forecasting (LTSF) due to their linear complexity. S4 models use four key components: $\Delta$ (discretization step size), $A$ (state update matrix), $B$ (input matrix), and $C$ (output matrix), operating in two modes: linear recurrence for autoregressive inference and global convolution for parallelizable training. The recurrent mode stores historical information similar to RNNs, while the convolutional mode transforms the recurrence into a convolution via Toeplitz matrices, making computation more efficient. However, S4 struggles with time-invariance issues, limiting its ability to selectively memorize or forget data. Mamba  [7] addresses this by making $B$, $C$, and $\Delta$ dynamic functions of the input, creating time-dependent parameters that adapt to varying contexts, filtering noise and maintaining Transformer-level performance with linear complexity. SIMBA [23] further enhances these models by incorporating a simplified block-sparse attention mechanism, blending state space models with attention to efficiently capture long-range dependencies while reducing computational overhead. This hybrid approach allows SIMBA to handle large-scale, noisy, or irregular time series more effectively than purely recurrent models like S4. TimeMachine[1] extends these innovations by using a quadruple Mamba setup to manage both channel mixing and channel independence. TimeMachine avoids the quadratic complexity of Transformers like iTransformer and PatchTST by using multi-scale context generation and selective scanning, enabling it to capture long-range dependencies efficiently while maintaining high performance in long-term forecasting tasks.
@@
RWKV-TS [11], is a novel linear RNN architecture specifically designed for time series tasks. Compared to traditional RNN models like LSTM and GRU, RWKV-TS achieves O(L) time and memory complexity, enhances long-range information capturing capabilities, and exhibits high computational efficiency and scalability. Orvieto et al. (2023 [22]) demonstrated that carefully designed deep RNNs can match the performance of deep state-space models (SSMs) on long-range reasoning tasks. They introduced the Linear Recurrent Unit (LRU), an RNN block that achieves similar performance to S4/S4D/S5 on the Long Range Arena benchmark while maintaining their computational efficiency. TTT  [24] layers introduce a new approach where the hidden state is treated as a trainable machine learning model and it learns not only during training but at test time as well. Instead of using a single fixed hidden state vector, TTT employs dynamically updated weights (in the form of matrices inside linear or MLP layers) to compress and store historical data. This dynamic adjustment during test time allows TTT to better capture long-term relationships by continuously incorporating new information. TTT also utilizes self-supervised training during the testing process, updating model weights in real-time to reconstruct corrupted input tokens. This method provides an efficient alternative to self-attention, maintaining linear complexity while offering the same interface as RNN layers, making it easily integrable into larger architectures. Additionally, TTT is designed to parallelize effectively, overcoming the typical scaling issues of RNNs. By improving the ability to process longer contexts with linear complexity, TTT offers a compelling alternative to Transformer-based architectures, which struggle with quadratic complexity due to self-attention.
@@
Recent advancements in long-term time series forecasting (LTSF) have introduced more efficient architectures that avoid the complexity of attention mechanisms and recurrence. One notable model is TSMixer [2], an MLP-based approach that separates temporal and feature interactions through time-mixing and channel-mixing operations, achieving linear scaling with sequence length. Despite its simplicity, TSMixer performs competitively across benchmarks, showing that complex architectures like Transformers and RNNs are not always necessary. However, MLP-based models can struggle with capturing long-range dependencies, and their performance may depend heavily on hyperparameter tuning, especially for smaller or highly non-linear datasets. Convolutional neural networks (CNNs) have also proven effective for LTSF, particularly in capturing local temporal patterns. ModernTCN[5] revisits temporal convolution networks (TCNs) and introduces CNN-based architectures for efficient long-term forecasting. By using dilated convolutions and a hierarchical structure, ModernTCN captures both short- and long-range dependencies while maintaining computational efficiency, making it suitable for multi-scale time series data. Building on these developments, we improve the original TimeMachine model by replacing its Mamba blocks with Test-Time Training (TTT) blocks to enhance long-context prediction capabilities. We also explore CNN configurations, such as convolutional stacks, to enrich local temporal feature extraction. This hybrid approach combines the efficiency of MLPs, the local pattern recognition of CNNs, and the global context modeling of TTT, leading to a more robust architecture for LTSF tasks that balances both short- and long-term forecasting needs.
@@ 3 Model Architecture
Our goal is to improve upon the performance of the state-of-the-art (SOTA) models in LTSF using the latest advancements in sequential modeling. Specifically, we integrate Test-Time Training (TTT) modules  [24] into our model for two key reasons, TTT is theoretically proven to have an extremely long context window, being a form of linear RNN  [22], capable of capturing long-range dependencies efficiently. Secondly, the expressive hidden states of TTT allow the model to capture a diverse set of features without being constrained by the architecture, including the depth of the hidden layers, their size, or the types of blocks used.
@@ 3.1 General Architecture
Our model features a two-level hierarchical architecture that captures both high-resolution and low-resolution context, as illustrated in Figure 1. To adapts to the specific characteristics of the dataset, the architecture handles two scenarios—Channel Mixing and Channel Independence—illustrated in Figure 1 and 1 respectively, while Figure 1 depicts the Channel Mixing scenario. A more detailed and mathematical description of the normalization and prediction procedures can be found in Appendix B.
@@ 4 Experiments and Evaluation
To validate the new architecture, we conducted extensive experiments, exploring various configurations of model parameters to assess the impact on performance. This included evaluating different sequence and prediction lengths and experimenting with alternative architectures applied in cascade on the forecasting task, leveraging the flexibility provided by TTT modules.
@@ 4.1 Original Experimental Setup
We evaluate our model on seven benchmark datasets that are commonly used for LTSF, namely: Traffic, Weather, Electricity, ETTh1, ETTh2, ETTm1, and ETTm2 from Wu et al. [28], and Zhou et al. [31]. Among these, the Traffic and Electricity datasets are significantly larger, with 862 and 321 channels, respectively, and each containing tens of thousands of temporal points.
Table 1 summarizes the dataset details:
Dataset Channels Time Points Frequencies
Weather 21 52696 10 Minutes
Traffic 862 17544 Hourly
Electricity 321 26304 Hourly
ETTh1 7 17420 Hourly
ETTh2 7 17420 Hourly
ETTm1 7 69680 15 Minutes
ETTm2 7 69680 15 Minutes
Table 1: Details of each dataset
@@ 4.2 Quantitative Results
Table 2: Results in MSE and MAE (the lower the better) for the long-term forecasting task. We compare extensively with baselines under different prediction lengths, $T=\{96,192,336,720\}$ following the setting of iTransformer [17]. The length of the input sequence ($L$) is set to 96 for all baselines. The best results are in bold and the second best are underlined.
@@
In the Electricity dataset, TTT showed dominant results across all horizons, achieving an MSE of 0.140 at horizon 96 and 0.201 at horizon 720, outperforming TimeMachine and PatchTST. For ETTh1, TTT was highly competitive, with strong short-term results (MSE of 0.358 at horizon 96) and continued dominance at medium-term horizons like 336, with an MSE of 0.429. While ETTh2 showed TimeMachine slightly ahead at short horizons, TTT closed the gap at longer horizons (MSE of 0.445 at horizon 720 compared to 0.411 for TimeMachine).
For the ETTm1 dataset, TTT outperformed TimeMachine at nearly every horizon, recording an MSE of 0.315 at horizon 96 and 0.431 at horizon 720, confirming its effectiveness for long-term industrial forecasting. Similarly, in ETTm2, TTT consistently performed best at shorter horizons (MSE of 0.177 at horizon 96) and remained highly competitive at longer horizons, maintaining its lead over TimeMachine at horizon 720 (MSE of 0.362 vs. 0.371).
@@ 5.1 Experimental Setup with Enhanced Architectures
To assess the impact of enhancing the model architecture, we conducted experiments by adding hidden layer architectures before the sequence modeling block in each of the four TTT blocks. The goal was to improve performance by enriching feature extraction through local temporal context. As shown in Figure 3. We tested the following configurations: (1) Conv 3: 1D Convolution with kernel size 3, (2) Conv 5: 1D Convolution with kernel size 5, (3) Conv Stack 3: two 1D Convolutions with kernel size 3 in cascade, (4) Conv Stack 5: two 1D Convolutions with kernel sizes 5 and 3 in cascade, and (5) Inception: an Inception Block combining 1D Convolutions with kernel sizes 5 and 3, followed by concatenation and reduction to the original size. Kernel sizes were limited to 5 to avoid oversmoothing, and original data dimensions were preserved to ensure consistency with the TTT architecture.
@@
Our findings reveal that the introduction of additional hidden layer architectures, including convolutional layers, had varying degrees of impact on performance across different horizons. The best-performing setup was the Conv Stack 5 architecture, which achieved the lowest MSE and MAE at the 96 time horizon, with values of 0.261 and 0.289, respectively, outperforming the TimeMachine model at this horizon. At longer horizons, such as 336 and 720, Conv Stack 5 continued to show competitive performance, though the gap between it and the TimeMachine model narrowed. For example, at the 720 horizon, Conv Stack 5 showed an MAE of 0.373, while TimeMachine had an MAE of 0.378.
However, other architectures, such as Conv 3 and Conv 5, provided only marginal improvements over the baseline TTT architectures (Linear and MLP). While they performed better than Linear and MLP, they did not consistently outperform more complex setups like Conv Stack 3 and 5 across all horizons. This suggests that hidden layer expressiveness can enhance model performance and that deeper convolutional layers that are more complex show supremacy over shallower architectures. However, added complexity can lead to slower training and inference times and comparing the results from Conv 3 and Conv Stacked 3 and Conv Stacked 5 we can see that Conv 3 performs better than Conv Stacked 3 but not as good as Conv Stacked 5 hinting on the compromise between simplicity and performance especially in situations where resources are constrained.
Further exploration into more sophisticated architectures, such as varying convolutional depths or integrating more advanced attention mechanisms, may provide additional opportunities for performance improvements. The current findings indicate that while convolutional layers can be beneficial, more diverse or deeper hidden layers may offer greater potential in future iterations of the model.
@@ 5.3 Results and Statistical Comparisons for Proposed Architectures
Table 4: MSE and MAE performance metrics for TimeMachine, TTT, and different convolutional architectures across prediction horizons.
General Performance
The proposed architectures—TTT Linear, TTT MLP, Conv Stack 3, Conv Stack 5, Conv 3, Conv 5, and Inception—demonstrate varying levels of performance across prediction horizons. TTT Linear performs well at shorter horizons, with an MSE of 0.268 and MAE of 0.298 at horizon 96, but experiences increasing error at longer horizons, with an MSE of 0.357 at horizon 336. TTT MLP follows a similar trend but with slightly worse overall performance. Conv 3 and Conv 5 outperform the Linear and MLP models at shorter horizons, achieving comparable MSE (0.269) and MAE (0.297) at horizon 96, but fall behind the more complex Conv Stack models at longer horizons. Conv Stack 5 performs best at shorter horizons with an MSE of 0.261 and MAE of 0.289 at horizon 96, while Conv Stack 3 shows a slight advantage at longer horizons with an MSE of 0.359 at horizon 336. The Inception architecture provides stable performance across horizons, closely following the Conv Stack models, with an MSE of 0.361 at horizon 336.
Impact of Architectures
The Conv Stack 5 architecture demonstrates the best overall performance among all convolutional models, particularly at shorter horizons, achieving the lowest MSE (0.261) and MAE (0.289) at horizon 96. Conv 3 and Conv 5 perform better than the simpler Linear and MLP models but are consistently outperformed by the more complex stacked architectures. At horizon 720, Conv 5 shows a marginal improvement over Conv 3, with an MSE of 0.371 compared to 0.377. The Inception architecture performs similarly to Conv Stack 5, offering consistent results across all horizons and proving particularly effective for complex, long-term predictions. Across all horizons, MSE and MAE increase as the horizon lengthens, but Conv Stack 3, Conv Stack 5, and Inception handle this increase more effectively than TimeMachine and other simpler models, which exhibit approximately 5% to 10% higher errors at longer horizons.

### v2

@@ 1 Introduction
  • Hidden Layer Architectures: The ablation studies revealed that while convolutional architectures added to the TTT modules provided some improvements, Conv Stack 5 consistently delivered the best results among the convolutional variants. However, simpler architectures like Conv 3 often performed comparably, showing that increased architectural complexity did not always lead to significantly better performance. Very complex architectures like the modern convolutional block from [5] showed competitive performance when used as TTT hidden layer architectures compared to the simpler single architectures proposed, hinting on the potential of more complex architectures in capturing more long term dependencies.
  • Adaptability to Long-Term Predictions: The TTT-based models excelled in long-term prediction tasks, especially for really high prediction lengths like 2880. TTT-based models also excelled on increased sequence lengths as high as 5760 which is the maximum sequence length allowed by the benchmark datasets. This verified the theoretically expected superiority of TTT based models relative to the mamba/transformer based SOTA models.
@@
In this work, we explore the potential of TTT modules in Long-Term Series Forecasting (LTSF) by integrating them into novel model configurations to surpass the current state-of-the-art (SOTA) models. Our key contributions are as follows:
@@
  • We conduct an ablation study to assess the performance of various hidden layer architectures within our model. By testing six different convolutional configurations, one of which being ModernTCN by [5], we quantify their impact on model performance and provide insights into how they compare with the SOTA model.
@@ 2 Related Work
Several Transformer-based models have advanced long-term time series forecasting (LTSF), with notable examples like iTransformer [18] and PatchTST
 [21]. iTransformer employs multimodal interactive attention to capture both temporal and inter-modal dependencies, suitable for multivariate time series, though it incurs high computational costs when multimodal data interactions are minimal. PatchTST, inspired by Vision Transformers [6], splits input sequences into patches to capture dependencies effectively, but its performance hinges on selecting the appropriate patch size and may reduce model interpretability. Other influential models include Informer [31], which uses sparse self-attention to reduce complexity but may overlook finer details in multivariate data; Autoformer [28], which excels in periodic data but struggles with non-periodic patterns; Pyraformer [16], which captures multi-scale dependencies through a hierarchical structure but at the cost of increased computational requirements; and Fedformer [32], which combines time- and frequency-domain representations for efficiency but may underperform on noisy time series. While each model advances LTSF in unique ways, they also introduce specific trade-offs and limitations.
@@
S4 models [8, 9, 10] are efficient sequence models for long-term time series forecasting (LTSF), leveraging linear complexity through four key components: $\Delta$ (discretization step size), $A$ (state update matrix), $B$ (input matrix), and $C$ (output matrix). They operate in linear recurrence for autoregressive inference and global convolution for parallel training, efficiently transforming recurrences into convolutions. However, S4 struggles with time-invariance issues, limiting selective memory. Mamba [7] addresses this by making $B$, $C$, and $\Delta$ dynamic, creating adaptable parameters that improve noise filtering and maintain Transformer-level performance with linear complexity. SIMBA [23] enhances S4 by integrating block-sparse attention, blending state space and attention to efficiently capture long-range dependencies while reducing computational overhead, ideal for large-scale, noisy data. TimeMachine [1] builds on these advances by employing a quadruple Mamba setup, managing both channel mixing and independence while avoiding Transformers’ quadratic complexity through multi-scale context generation, thereby maintaining high performance in long-term forecasting tasks.
@@
RWKV-TS [11] is a novel linear RNN architecture designed for time series tasks, achieving O(L) time and memory complexity with improved long-range information capture, making it more efficient and scalable compared to traditional RNNs like LSTM and GRU. [22] introduced the Linear Recurrent Unit (LRU), an RNN block matching the performance of S4 models on long-range reasoning tasks while maintaining computational efficiency. TTT [24] layers take a novel approach by treating the hidden state as a trainable model, learning during both training and test time with dynamically updated weights. This allows TTT to capture long-term relationships more effectively through real-time updates, providing an efficient, parallelizable alternative to self-attention with linear complexity. TTT’s adaptability and efficiency make it a strong candidate for processing longer contexts, addressing the scalability challenges of RNNs and outperforming Transformer-based architectures in this regard.
@@
Recent advancements in long-term time series forecasting (LTSF) have introduced efficient architectures that avoid the complexity of attention mechanisms and recurrence. TSMixer [2], an MLP-based model, achieves competitive performance by separating temporal and feature interactions through time- and channel-mixing, enabling linear scaling with sequence length. However, MLPs may struggle with long-range dependencies and require careful hyperparameter tuning, especially for smaller datasets. Convolutional neural networks (CNNs) have also proven effective for LTSF, particularly in capturing local temporal patterns. ModernTCN [5] improves temporal convolution networks (TCNs) using dilated convolutions and a hierarchical structure to efficiently capture both short- and long-range dependencies, making it well-suited for multi-scale time series data.
Building on these developments, we improve the original TimeMachine model by replacing its Mamba blocks with Test-Time Training (TTT) blocks to enhance long-context prediction capabilities. We also explore CNN configurations, such as convolutional stacks, to enrich local temporal feature extraction. This hybrid approach combines the efficiency of MLPs, the local pattern recognition of CNNs, and the global context modeling of TTT, leading to a more robust architecture for LTSF tasks that balances both short- and long-term forecasting needs.
@@ 3 Model Architecture
Our goal is to improve upon the performance of the state-of-the-art (SOTA) models in LTSF using the latest advancements in sequential modeling. Specifically, we integrate Test-Time Training (TTT) modules into our model for two key reasons, TTT is theoretically proven to have an extremely long context window, being a form of linear RNN [22], capable of capturing long-range dependencies efficiently. Secondly, the expressive hidden states of TTT allow the model to capture a diverse set of features without being constrained by the architecture, including the depth of the hidden layers, their size, or the types of blocks used.
@@ 3.1 General Architecture
Our model features a two-level hierarchical architecture that captures both high-resolution and low-resolution context, as illustrated in Figure 1. To adapts to the specific characteristics of the dataset, the architecture handles two scenarios—Channel Mixing and Channel Independence—illustrated in Figure 1 and 1 respectively. A more detailed and mathematical description of the normalization and prediction procedures can be found in Appendix B.
@@ 4.1 Original Experimental Setup
We evaluate our model on seven benchmark datasets that are commonly used for LTSF, namely: Traffic, Weather, Electricity, ETTh1, ETTh2, ETTm1, and ETTm2 from  [28] and  [31]. Among these, the Traffic and Electricity datasets are significantly larger, with 862 and 321 channels, respectively, and each containing tens of thousands of temporal points. Table 6 summarizes the dataset details in Appendix D.
@@ 4.2 Quantitative Results
Table 1: Results in MSE and MAE (the lower the better) for the long-term forecasting task. We compare extensively with baselines under different prediction lengths, $T=\{96,192,336,720\}$ following the setting of iTransformer [17]. The length of the input sequence ($L$) is set to 96 for all baselines. TTT (ours) is our TTT block with the Conv Stack 5 architecture. The best results are in bold and the second best are underlined.
@@
In the Electricity dataset, TTT showed dominant results across all horizons, achieving an MSE of 0.140 at horizon 96 and 0.201 at horizon 720, outperforming TimeMachine and PatchTST. For ETTh1, TTT was highly competitive, with strong short-term results (MSE of 0.358 at horizon 96) and continued dominance at medium-term horizons like 192, with an MSE of 0.409. While ETTh2 showed TimeMachine slightly ahead at short horizons, TTT closed the gap at longer horizons (MSE of 0.445 at horizon 720 compared to 0.411 for TimeMachine).
For the ETTm1 dataset, TTT outperformed TimeMachine at nearly every horizon, recording an MSE of 0.315 at horizon 96 and 0.431 at horizon 720, confirming its effectiveness for long-term industrial forecasting. Similarly, in ETTm2, TTT remained highly competitive at longer horizons, with a lead over TimeMachine at horizon 720 (MSE of 0.362 vs. 0.371). The radar plot in Figure 2 shows the comparison between TTT (ours) and TimeMachine for both MSE and MAE on all datasets.
@@ 5.1 Experimental Setup with Enhanced Architectures
To assess the impact of enhancing the model architecture, we conducted experiments by adding hidden layer architectures before the sequence modeling block in each of the four TTT blocks. The goal was to improve performance by enriching feature extraction through local temporal context. As shown in Figure 3 in Appendix B.
We tested the following configurations: (1) Conv 3: 1D Convolution with kernel size 3, (2) Conv 5: 1D Convolution with kernel size 5, (3) Conv Stack 3: two 1D Convolutions with kernel size 3 in cascade, (4) Conv Stack 5: two 1D Convolutions with kernel sizes 5 and 3 in cascade, (5) Inception: an Inception Block combining 1D Convolutions with kernel sizes 5 and 3, followed by concatenation and reduction to the original size and (6) ModernTCN: A modern convolutional block proposed in [5] that uses depthwise and pointwise convolutions with residual connections similar to the structure of the transformer block. For the simpler architectures kernel sizes were limited to 5 to avoid oversmoothing, and original data dimensions were preserved to ensure consistency with the TTT architecture. For ModernTCN we reduced the internal dimensions to 16 (down from the suggested 64) and did not use multiscale due to the exponential increase in GPU memory required which slowed down the training process and did not allow the model to fit in a single A100 GPU. We kept the rest of the parameters of ModernTCN the same as in the original paper.
@@
Our findings reveal that the introduction of additional hidden layer architectures, including convolutional layers, had varying degrees of impact on performance across different horizons. The best-performing setup was the Conv Stack 5 architecture, which achieved the lowest MSE and MAE at the 96 time horizon, with values of 0.261 and 0.289, respectively, outperforming the TimeMachine model at this horizon. At longer horizons, such as 336 and 720, Conv Stack 5 continued to show competitive performance, with a narrow gap between it and the TimeMachine model. For example, at the 720 horizon, Conv Stack 5 showed an MAE of 0.373, while TimeMachine had an MAE of 0.378.
However, other architectures, such as Conv 3 and Conv 5, provided only marginal improvements over the baseline TTT architectures (Linear and MLP). While they performed better than Linear and MLP, they did not consistently outperform more complex setups like Conv Stack 3 and 5 across all horizons. This suggests that hidden layer expressiveness can enhance model performance.
ModernTCN showed competitive results across multiple datasets (see Appendix D), such as ETTh2, where it achieved an MSE of 0.285 at horizon 96, outperforming Conv 3 and Conv 5. However, as with other deep convolutional layers, ModernTCN’s increased complexity also led to slower training times compared to simpler setups like Conv 3 and it failed to match Conv Stack 5’s performance.
@@ 5.3 Results and Statistical Comparisons for Proposed Architectures
Table 3: MSE and MAE performance metrics for TimeMachine, TTT blocks with original architectures (MLP & Linear), and TTT block with different convolutional architectures across all prediction horizons.
The proposed architectures—TTT Linear, TTT MLP, Conv Stack 3, Conv Stack 5, Conv 3, Conv 5, Inception, and TTT with ModernTCN—exhibit varying performance across prediction horizons. TTT Linear performs well at shorter horizons (MSE 0.268, MAE 0.298 at horizon 96) but declines at longer horizons (MSE 0.357 at horizon 336). TTT MLP follows a similar trend with slightly worse performance. Conv 3 and Conv 5 outperform Linear and MLP at shorter horizons (MSE 0.269, MAE 0.297 at horizon 96) but lag behind Conv Stack models at longer horizons. TTT with ModernTCN shows promising results at shorter horizons, such as MSE 0.389, MAE 0.402 on ETTh1, and MSE 0.285, MAE 0.340 on ETTh2 at horizon 96. Although results for Traffic and Electricity datasets are pending, preliminary findings indicate TTT with ModernTCN is competitive, particularly for short-term dependencies (see Table 7 in Appendix D). Conv Stack 5 performs best at shorter horizons (MSE 0.261, MAE 0.289 at horizon 96). Inception provides stable performance across horizons, closely following Conv Stack 3 (MSE 0.361 at horizon 336). At horizon 720, Conv 5 shows a marginal improvement over Conv 3, with an MSE of 0.400 compared to 0.406. The Conv Stack 5 architecture demonstrates the best overall performance among all convolutional models.